-5 1/2 + (-2 1/2)=
Answer by checkley79(3341)

-5 1/2 + (-2 1/2)
-(5*2+1)/2-(2*2+1)/2
-(10+1)/2-(4+1)/2
-11/2-5/2=-16/2=-8 ans.